Plastic drum of cylindrical shape is made by melting spherical solid plastic balls of radius 1cm .Find the number of balls required to make the drum of thickness 2cm,height 90cm and outer radius 30cm
Found 2 solutions by Alan3354, Sanchita:
Answer by Alan3354(69443)

Plastic drum of cylindrical shape is made by melting spherical solid plastic balls of radius 1cm. Find the number of balls required to make the drum of thickness 2cm, height 90cm and outer radius 30cm
-----------------
Does that include one end, 2 ends, no ends?
-----------
With one end closed:
Vol = pi*h*(30^2 - 28^2) + pi*30^2*2
Vol = pi*88*116 + 1800*pi
Vol = 12008*pi cc
------
Divide by the volume of 1 ball.
--> 16011 balls

Answer by Sanchita(1)

Vol.of cylinder=πr2h
=(22/7)*(30^2)*90
=254571.43
Vol.of inner cylinder=πr2h
=(22/7)*(28^7)*88
=216832
Vol.of cylinder=vol.of outer cylinder-vol.of inner cylinder
=254571.43-216832
=37739.43 cm^3
Vol.of 1 ball=(4/3)*(22/7)*1^3
4.1905
No.of ball required =37739.43/4.1905
=9006 balls